Imagine making waves on a slinky. First, we make a wave with: wavelength 2 cm long, amplitude 0.2 cm, Frequency at 1 Hz and wave speed of 2 cm/s.

If we then shake the slinky twice as often, what happens to the wavelength, amplitude, frequency and wave speed?
1
Expert's answer
2019-04-29T09:29:52-0400

Frequency will be doubled:


"f'=2f=2(1)=2\ Hz"


The wave speed will be the same:


"v'=v= 2\frac{cm}{s}"

The wavelength will be:


"\lambda'=\frac{v}{f'}=\frac{2}{2}=1\ cm = \frac{1}{2} \lambda"

The amplitude will be the same:


"A'=A=0.2 \ cm"